The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

          mcnews@astro.uiuc.edu
http://www.astro.uiuc.edu/mcnews/MCNews.html
  http://www.astro.uni-bonn.de/~mcnews/

No. 22 July 5, 1998

Contents

News  1
Abstracts of 7 refereed papers  2
Abstracts of 2 non-refereed papers 6
Abstract of a book  7

News

OGLE data available over the Internet

The Optical Gravitational Lensing Experiment (OGLE) team announces the availability of:

• “OGLE BV I maps of Dense Stellar Regions. The Small Magellanic Cloud.” The maps contain BV I photometric and astrometric data of about 2.2 million stars from the central regions of the SMC (about 2.4 square degrees) observed during the OGLE microlensing survey.

• “Catalog of Star Clusters in the Small Magellanic Cloud.” The catalog contains data and BV I photometry of 238 star clusters from the central part of the SMC.

The maps and catalog can be retrieved from the following URLs:

http://www.astrouw.edu.pl/~ftp/ogle or its mirror http://www.astro.princeton.edu/~ogle

Andrzej Udalski

Abstracts of Refereed Papers

Compact HII regions in the Large Magellanic Cloud observed by ISO

F. Comerón ^{1} and P. Claes^{2}

1 European Southern Observatory, Karl-Schwarzschild-Strasse 2, 85748 Garching, Germany 2 ESTEC, European Space Agency, NL 2200 AG Noordwijk, The Netherlands

We present ISOCAM observations of five fields in the Large Magellanic Cloud centered on IRAS point sources with colors of ultracompact HII regions. Three bright sources of parsec or subparsec size, with luminosities consistent with being embedded O stars, are detected at 15 μm in two fields belonging to the HII region N 159. Using published high resolution radio observations, it is found that two of these sources may be ultracompact HII regions with a turnover frequency above 2.4 GHz, while the other is identified with a known compact HII region. If their nature as true ultracompact HII regions is confirmed by follow-up observations, these objects can provide a first sample useful for future studies of early massive star evolution in a low metallicity environment.

Accepted by: Astronomy and Astrophysics (Letters) For preprints, contact fcomeron@eso.org

Also available from the URL http://www.eso.org/∼fcomeron/publicat.html

Star clusters and the structure of the ISM. Tunnels and wakes in giant extragalactic HII regions

G. Tenorio-Tagle^{1} and G. Medina Tanco^{2,3}

1 Instituto Nacional de Astrofísica Optica y Electrónica, AP 51, 72000 Puebla, México 2 Instituto Astronômico e Geofísico, Universidade de São Paulo, Brazil; gustavo@iagusp.usp.br 3 Dept. of Physics and Astronomy, University of Leeds, Leeds LS2 9JT, UK

Several structures have been discovered embedded in regions of recent or ongoing star formation, which point to the importance of the interaction between fast moving wind-blowing stars and their environment. Using hydrodynamic simulations, we investigate the passage through the interstellar medium of a supersonic stellar wind source, and show how it can naturally lead to the formation of tubes, channels and swamps of globules as interfaces are crossed. The results are in excellent agreement with observation of 30 Doradus.

Accepted by: Astrophysical Journal Letters For preprints, contact Gustavo Medina Tanco, gustavo@iagusp.usp.br

Also available from the URL http://babbage.sissa.it/abs/astro-ph/9806367

HST UV measurements of wind structure and velocities in Local Group OB stars

Raman K. Prinja and Paul A. Crowther

Department of Physics & Astronomy, University College London, Gower Street, London WC1E 6BT, UK

Archival HST FOS and GHRS data sets have been used to collect ultraviolet evidence for largeand small-scale stellar wind structure in extragalactic Local Group OB stars (i.e. SMC, LMC including R136, M 31, M 33, and NGC 6822). By comparison to previous studies of Galactic OB stars, wind activity is principally diagnosed in individual spectrograms via the presence of ‘narrow absorption components’ and saturated ‘black’ absorption troughs in the resonance line doublets. Their characteristics broadly suggest that these stars share the same physical mechanisms for perturbing the winds as those that act in Galactic stars. Both these spectral indicators are also used to provide reliable measures of wind terminal velocities. These velocities are directly compared to previously published Galactic values, without reliance on model profile fitting. Relative to Galactic OB stars, the most discrepant terminal velocities (and wind line profiles) are due to main-sequence early O-type stars in the SMC.

Accepted by: Monthly Notices of the Royal Astronomical Society For preprints, contact rkp@star.ucl.ac.uk

Discovery of H_{2}O maser emission from the red supergiant IRAS04553−6825 in the Large Magellanic Cloud

Jacco Th. van Loon^{1}, Peter te Lintel Hekkert^{2}, Valentín Bujarrabal^{3}, Albert A. Zijlstra^{4} and Lars-Åke Nyman^{5,6}

1 Astronomical Institute “Anton Pannekoek”, University of Amsterdam, Kruislaan 403, NL-1098 SJ Amsterdam, The Netherlands 2 Australia Telescope National Facility, Parkes Observatory, P.O.Box 276, Parkes, NSW 2870, Australia 3 Observatorio Astronómico Nacional, Campus Universitario, Apartado 1143, E-28800 Alcalá de Henares (Madrid), Spain 4 University of Manchester Institute of Science and Technology, P.O.Box 88, Manchester M60 1QD, United Kingdom 5 European Southern Observatory, Casilla 19001, Santiago 19, Chile 6 Onsala Space Observatory, S-439 92 Onsala, Sweden

We report the detection of 22 GHz H_{2}O maser emission from the red supergiant IRAS04553−6825 in the Large Magellanic Cloud. It is the first known source of circumstellar H_{2}O maser emission outside the Milky Way. The measured flux density is comparable to that expected from scaling the galactic red supergiant NML Cyg. The peak velocity agrees with the SiO maser peak velocity.

A near-infrared spectrum indicates that IRAS04553−6825 has a typical LMC metallicity. We argue that, possibly as a result of the low metallicity, the H_{2}O emission probably occurs near or within the dust formation radius, rather than further out as appears to be the case in NML Cyg and galactic OH/IR stars.

Accepted by: Astronomy & Astrophysics

For preprints, contact jvanloon@astro.uva.nl

Hubble Space Telescope Images of Magellanic Cloud Planetary Nebulae: Data and Correlations across Morphological Classes

L. Stanghellini^{1,2,3}, J.C. Blades^{1}, S.J. Osmer^{1,4}, M.J. Barlow^{5}, and X.–W. Liu^{5}

1 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore MD 21218, USA 2 Affiliated to the Astrophysics Division, Space Science Department of ESA 3 Osservatorio Astronomico di Bologna, Italy 4 Present address: Department of Computer Sciences, University of Edinburgh, King’s Buildings, Mayfield Road, Edinburgh EH9 3JZ, Scotland 5 Department of Physics and Astronomy, University College London, Gower Street, London WC1 6BT, UK

The morphology of planetary nebulae (PNe) provides an essential tool for understanding their origin and evolution, as it reflects both the dynamics of the gas ejected during the TP–AGB phase, and the central star energetics. Here we study the morphology of 27 Magellanic Cloud planetary nebulae (MCPNe) and present an analysis of their physical characteristics across morphological classes. Similar studies have been successfully carried out for galactic PNe, but were compromised by the uncertainty of individual PN distances. We present our own HST/FOC images of 15 Magellanic Cloud PNe (MCPNe) acquired through a narrow-band λ5007 [O iii] filter. We use the Richardson–Lucy deconvolution technique on these pre–COSTAR images to achieve post–COSTAR quality. Three PNe imaged before and after COSTAR confirm the high reliability of our deconvolution procedure. We derive morphological classes, dimensions, and surface photometry for all these PNe. We have combined this sample with HST/PC1 images of 15 MCPNe, three of which are in common with the FOC set, acquired by Dopita et al. (1996), to obtain the largest MCPN sample ever examined from the morphological viewpoint. By using the whole database, supplemented with published data from the literature, we have analyzed the properties of the MCPNe and compared them to a typical, complete galactic sample. Morphology of the MCPNe is then correlated with PN density, chemistry, and evolution.

Accepted by: The Astrophysical Journal For preprints, contact lstanghe@stsci.edu

The Optical Gravitational Lensing Experiment. BVI Maps of Dense Stellar Regions. I. The Small Magellanic Cloud.

A. Udalski^{1}, M. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1}, P. Woźniak^{2}, and K. Żebruń^{1}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2 Princeton University Observatory, Princeton, NJ 08544-1001, USA

We present three color, BVI maps of the Small Magellanic Cloud. The maps contain precise photometric and astrometric data for about 2.2 million stars from the central regions of the SMC bar covering ≈ 2.4 square degrees on the sky. Mean brightness of stars is derived from observations collected in the course of the OGLE-II microlensing search from about 130, 30 and 15 measurements in the I, V and B-bands, respectively. Accuracy of the zero points of photometry is about 0.01 mag, and astrometry 0.15 arcsec (with possible systematic error up to 0.7 arcsec). Color-magnitude diagrams of observed fields are also presented.

The maps of the SMC are the first from the series of similar maps covering other OGLE fields: LMC, Galactic bulge and Galactic disk. The data are very well suited for many projects, particularly for the SMC which has been neglected photometrically for years. Because of potentially great impact on many astrophysical fields we decided to make the SMC data available to the astronomical community from the OGLE Internet archive.

Accepted by: Acta Astronomica 1998, 48, 147

For preprints, contact udalski@sirius.astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9806313,
http://www.astrouw.edu.pl/~ftp/ogle, and http://www.astro.princeton.edu/~ogle

The Optical Gravitational Lensing Experiment. The Catalog of Clusters in the Small Magellanic Cloud.

G. Pietrzyński^{1}, A. Udalski^{1}, M. Kubiak^{1}, M. Szymański^{1}, P. Woźniak^{2}, and K. Żebruń^{1}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2 Princeton University Observatory, Princeton, NJ 08544-1001, USA

We present the catalog of clusters found in the area of ≈ 2.4 square degree in the central regions of the Small Magellanic Cloud. The catalog contains data for 238 clusters, 72 of them are new objects. For each cluster equatorial coordinates, radii, approximate number of members, cross-identification, finding chart and color magnitude diagrams: V − (B − V) and V − (V − I) are provided. Photometric data for all clusters presented in the catalog are available from the OGLE Internet archive.

Accepted by: Acta Astronomica 1998, 48, 175

For preprints, contact pietrzyn@sirius.astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9806321,
http://www.astrouw.edu.pl/~ftp/ogle, and http://www.princeton.edu/~ogle

Abstracts of Non-Refereed Papers

Hot Gas in the Large Magellanic Cloud

You-Hua Chu

Astronomy Department, University of Illinois, 1002 W. Green Street, Urbana, IL 61801, USA

Massive stars interact with the interstellar medium (ISM) via UV radiation, fast stellar winds, and supernova ejecta. These energetic interactions ionize, accelerate, and heat the ISM, producing hot gas with temperatures up to 10^{7} K. The mass exchange between this hot gas and the cold neutral gas constitutes the evolution of a multi-phase ISM.

Hot ionized gas is an essential component of the ISM. However, it is difficult to study the distribution and physical properties of hot gas in the Galaxy because of confusion and obscuration along the Galactic plane. The Large Magellanic Cloud (LMC), owing to its close proximity, nearly face-on view, and small extinction, presents an ideal site to study hot gas and its role in the interstellar matrix.

Soft X-ray images of the LMC show bright diffuse emission, indicating the existence of hot (10^{6}–10^{7} K) ionized gas. Clues to the origin of the hot gas in the LMC are provided by comparisons between X-ray and Hα images. Hα images of the LMC show a variety of interstellar shells, ranging from small supernova remnants a few pc in diameter to large supergiant shells more than 1000 pc across. Some of the diffuse X-ray emission is associated with shell structures, but there is also large-scale diffuse emission that does not appear to be confined by any cooler interstellar structures. The origins of hot gas in these different types of diffuse X-ray emission regions are individually discussed.

A hot (10^{5} K) gas halo of the LMC was suggested in the early 1980’s. As most of the previously reported interstellar C iv absorption occurs in large shell structures where supernovae and fast stellar winds are rampant, the evidence for a global hot gas halo around the LMC has been weak. Recent HST GHRS observations of five carefully chosen probe stars show clear detections of the interstellar C iv, providing unambiguous evidence for a hot gas halo of the LMC. Three sightlines have C iv velocities blue shifted from the Hα by 25–60 km s^{−1}, indicating that the hot gas halo may be expanding away from the disk gas. The hot gas halo of the LMC is most likely patchy, as implied by a stringent nondetection by the IUE along one sightline.

To appear in: Proceedings of the Workshop “The Large Magellanic Cloud and Other Dwarf Galaxies”, Bad Honnef, Jan 18-22, 1998, eds. T. Richtler & J.M. Braun For preprints, contact chu@astro.uiuc.edu

Also available from the URL http://www.astro.uiuc.edu/∼chu/publications/preprints/lmchg.ps

Star Formation Histories of Local Group Dwarf Galaxies

Eva K. Grebel

UCO/Lick Observatory, University of California at Santa Cruz, Santa Cruz, CA 95064, USA

40 galaxies are currently known as, or suspected to be Local Group members within a radius of 1.8 Mpc. 37 of these galaxies are dwarf galaxies with M_{B} > −18.5 mag. I present a compilation of the star formation histories of dwarf irregulars, dwarf ellipticals, and dwarf spheroidals in the Local Group, and visualize their evolutionary histories through Hodge’s population boxes.

The study of these star formation histories is a multi-parameter problem: Ages, metallicities, population fractions, and spatial variations must be determined, which depend crucially on the knowledge of reddening and distance.

Even the lowest-mass dwarf spheroidal galaxies can show significant spatial variations in star formation, ages, and metallicities. All Local Group dwarf galaxies appear to contain old (> 10 Gyr) and intermediate-age (1 − 10 Gyr) populations irrespective of morphological type. No two galaxies are alike. They vary widely in star formation histories, metallicities, fractions, and ages of their subpopulations even within the same morphological type. Star formation has occurred either in distinct episodes or continuously over long periods of time.

Invited review to be published in the proceedings of the XVIIIth Moriond conference on “Dwarf Galaxies and Cosmology”, Les Arcs, March 1998, eds. T.X. Thuan, C. Balkowski, V. Cayatte, & J. Tran Thanh Van, Editions Frontières For preprints, contact grebel@ucolick.org

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9806191

Book Abstract

The Brightest Binaries

D. Vanbeveren, W. Van Rensbergen, and C. De Loore

Astrophysical Institute, Vrije Universiteit Brussel, Pleinlaan 2, 1050 Brussel, Belgium

The monograph describes in extenso the present state of massive star research seen from the standpoint of stellar evolution, with special emphasis on close binaries. Statistics of massive close binaries are reasonably complete for the Solar neighbourhood. We demonstrate that within our knowledge, many scientific results where the effect of binaries are ignored, have an academic value but may be far from reality.

In chapter I, we summarize general observations of massive stars where we focus on the HR diagram, stellar wind mass loss rates, the stellar surface chemistry, rotation, circumstellar environments, supernovae.

Close binaries can not be studied separately from single stars and vice versa. First, the evolution of single stars is discussed (chapter II). We present new calculations with updated stellar wind mass loss rate formalisms. Especially the rates during the red supergiant phase and the rates during the Wolf-Rayet (WR) phase are different compared to previous values and this affects in a profound way the evolution during core helium burning. The computations are compared in detail to representative observations. The chapter ends with a general discussion on the effects of rotation on massive single star evolution.

Massive binaries, the main topic of the work, are considered in chapter III. Basic processes are described, i.e. the Roche lobe overflow and mass transfer, the common envelope process, the spiral-in process in binaries with extreme mass ratio, the mass accretion process and the merging process, the effect of the (asymmetric) supernova explosion of one of the components on the orbital parameters of the binary. Evolutionary computations of interacting close binaries are discussed and general conclusions are drawn. The enormous amount of observational data of massive binaries is summarized. We separately consider the non-evolved and evolved systems. The latter class includes the semidetached and contact binaries, the WR binaries, the X-ray binaries, the runaways, the single and binary pulsars. A general comparison between theoretical evolution and observations is combined with a discussion of specially interesting binaries: the evolved binaries HD 163181, HD 12323, HD 14633, HD 193516, HD 25638, HD 209481, ϕ Per and υ Sgr; the WR+OB binaries γ^{2}Vel, V444 Cyg; the high mass X-ray binaries Vela X-1, Wray 977, Cyg X-1; the low mass X-ray binaries Her X-1 and those with a black hole candidate; the runaways ζ Oph and ζ Pup, the WR+compact companion candidates HD 50896 and Cyg X-3; the supernova event SN 1987A. We finally propose an overall evolutionary model of massive close binaries as a function of primary mass, mass ratio and orbital period.

The fourth chapter deals with massive star population synthesis with a realistic population of binaries. We discuss the massive close binary frequency, mass ratio and period distribution, the observations that allow to constrain possible asymmetries during the supernova explosion of a massive star. Observed star numbers (as a function of metallicity) are then compared to theoretically predicted numbers where we focus on stellar populations in regions of continuous star formation and in starburst regions. Special attention is given to the O-type star/WR star/red supergiant star population, the pulsar and binary pulsar population, the supernova rates.

The monograph counts 347 pages, contains a considerable amount of new results and will be useful as reference work for everybody working in the field of massive stars in general, in the field of massive binaries in particular. Our hope is that this book encourages youngish researchers to pursue binary research, rather than fall into the bottomless pit of extra-Galactic research, including the eternal and largely illusory quest for ’standard candles’. Probably the best standard candles are, or will be, massive OB eclipsing systems, which in principle are capable of giving distances independent of the usual chain of hypotheses.

The book is published by Kluwer (Dordrecht), ISBN 0-7923-5155-X and can be ordered on-line via www.wkap.nl.